Solve
8y^2+xy=14
14y+x=25
Looking for order pairs.

Answer by Alan3354(69443)

8y^2+xy=14
14y+x=25 --> x = 25 - 14y
-------------
8y^2+xy=14
Sub for x
8y^2+ y*(25 - 14y) = 14
8y^2 + 25y - 14y^2 = 14
6y^2 - 25y + 14 = 0
(3y - 2)*(2y - 7) = 0
-------------
etc